# Cam4DOcc: Benchmark for Camera-Only 4D Occupancy Forecasting in Autonomous Driving Applications

### v2

@@ III-A Task Definition
Given $N_{p}$ past and the current consecutive camera images =$\mathcal{I}=\{I_{t}\}_{t=-N_{p}}^{0}$ as input, 4D occupancy forecasting aims to output the current occupancy $\mathbf{O}_{c}\in\mathbb{R}^{1\times H\times W\times L}$ and the future occupancy $\mathbf{O}_{f}\in\mathbb{R}^{N_{f}\times H\times W\times L}$ in a short time interval $N_{f}$, where $H$, $W$, $L$ represent the height, width, and length of the specific range defined in the present coordinate system ($t=0$). Each voxel of $\mathbf{O}_{f}$ has $N_{f}$ sequential states $\mathcal{S}=\{S_{t}\}_{t=1}^{N_{f}}$ to represent whether it is free or occupied in each future timestamp.
@@ III-B Dataset in New Format
Our Cam4DOcc benchmark introduces a new dataset format based on original nuScenes [29], nuScenes-Occupancy [13], and Lyft-Level5. As Fig. 2 illustrates, we first split the original nuScenes dataset into sequences with the time length of $N=N_{p}+N_{f}+1$. Then sequential semantic and instance annotations of movable objects are extracted for each sequence and collected into the GMO class, including bicycle, bus, car, construction, motorcycle, trailer, truck, and pedestrian. They are all transformed to the present coordinate system ($t=0$). After that, we voxelize the present 3D space and attach semantic/instance labels to the grids of movable objects using bounding boxes annotation. Notably, the invalid instance is discarded in this process once: (1) its visibility is under 40% over the 6 camera images if it is a newly appeared object in $N_{p}$ historical frames, (2) it first appears in $N_{f}$ incoming frames, or (3) it moves beyond the range ($H,W,L$) predefined at $t=0$. The visibility is quantified by the visible proportion of all pixels of the instance showing in camera images [29]. The sequential annotations are exploited to fill in missing intermediate instances based on constant velocity assumption [22, 44]. The same operations are also applied to the Lyft-Level5 dataset. The distribution of instance duration $[t_{in},t_{out}]$ after the processing mentioned above is presented in supplementary Sec. A. Lastly, we generate 3D backward centripetal flow using the instance association in the annotations. Li et al. [22] introduced 2D backward centripetal flow to improve the efficiency of 2D instance prediction. Inspired by that, we calculate 3D backward centripetal flow pointing from the voxel at time $t$ to its corresponding 3D instance center at $t-1$. We exploit this 3D flow to improve the accuracy of camera-based 4D occupancy forecasting (see Sec. V-C).
@@ V-A Experimental Setups
Baseline setups. We choose the state-of-the-art camera-based approaches as the outset of each baseline proposed in Sec. III-D. For the static-world occupancy model, we use the camera baseline of OpenOccupancy [13] (OpenOccupancy-C) to estimate the occupancy state of the present frame, which is then regarded as the prediction of all the future time steps. For the voxelization of point cloud prediction, we use SurroundDepth [50] to estimate continuous surrounding depth maps, which are then downsampled to generate pseudo point clouds by ray casting. Based on sequential pseudo point clouds input, we then use PCPNet [37] to forecast incoming 3D point clouds, followed by Cylinder3D [46] to extract point-level GMO and GSO labels, and further voxelize the results into occupancy grids (SPC). For the 2D-3D instance-based prediction, we choose PowerBEV [22] to forecast occupancy semantics on BEV and then lift the 2D results to 3D space (PowerBEV-3D). As to our proposed OCFNet, we directly implement 4D occupancy forecasting end-to-end. Notably, PowerBEV is trained by the 2D ground-truth semantics and 2D flow projected to the BEV plane. Besides, only PowerBEV and OCFNet are trained with flow annotations from Cam4DOcc simultaneously since they both have the flow head. To show that our proposed OCFNet can generate good forecasted results even seeing limited training data, we report the performance of OCFNet only trained on $\frac{1}{6}$ training sequences as well as the performance of the one trained on all training sequences (${}^{{\dagger}}$). OpenOccupancy-C, PowerBEV, and OCFNet are trained for 15 epochs using AdamW optimizer [51] with an initial learning rate 3e-4 and a weight decay of 0.01. SurroundDepth and Cylinder3D used in the point cloud prediction baseline are fine-tuned as their open sources suggest. PCPNet is firstly pretrained by range loss for 40 epochs using the same optimizer, but the initial learning rate is set to 1e-3. After that, it is further fine-tuned by Chamfer distance loss [52] for 10 epochs with a learning rate of 6e-4. All the networks mentioned above are trained with a batch size of 8 on 8 A100 GPUs. More details about the model parameters of our OCFNet are provided in supplementary Sec. B.
@@
Evaluation on forecasting inflated GMO. Results of the first task, forecasting inflated GMO on nuScenes and Lyft-Level5, are presented in Tab. I. Here, OpenOccupancy-C, PowerBEV, and OCFNet are trained only with inflated GMO labels, while PCPNet is trained by holistic point clouds. As shown, OCFNet and ${}^{{\dagger}}$ outperform all other baselines, surpassing the BEV-based method by 12.4% and 13.3% in $\text{IoU}_{f}$ and $\tilde{\text{IoU}}_{f}$ on nuScenes. On Lyft-Level5, our OCFNet and ${}^{{\dagger}}$ consistently outperforms PowerBEV-3D by 20.8% and 21.8% in $\text{IoU}_{f}$ and $\tilde{\text{IoU}}_{f}$. In addition, Fig. 5 shows the results of nuScenes GMO occupancy forecasted by our OCFNet and ${}^{{\dagger}}$, which indicates that OCFNet trained only with limited data can still capture the motion of GMO occupancy grids reasonably. The visualization on Lyft-Level5 is shown in supplementary Sec. F. The baseline SPC cannot work well for the present frame and even tends to fail while forecasting future occupancy state. This is because movable objects are labeled as the inflated dense voxel grids in this task, while the voxelization of PCPNet outputs is from sparse point-level prediction. In addition, the shape of the predicted objects loses consistency significantly in future time steps. The performance of OpenOccupancy-C is much better than that of the point cloud prediction baseline but still has a weak ability to estimate present occupancy and forecast future occupancy compared to PowerEBV-3D and OCFNet.
@@
Evaluation on forecasting fine-grained GMO. We further report the occupancy estimation and forecasting performance on fine-grained general movable objects with nuScenes-Occupancy (the second-level task). In Tab. II, we exhibit how the IoU of the forecasted objects changes once the GMO annotations have fine-grained voxel format rather than the inflated one in the first-level task for training as well as evaluation. It can be seen that the IoU of GMO forecasted by all the methods except the point cloud prediction baseline decreases significantly because it is rather difficult to predict sophisticated moving 3D structures using past continuous camera images. In contrast, SPC presents slightly better performance compared to the results in Tab. I since the ground-truth labels are also fine-grained and sparser than the counterparts in the first-level task. However, due to the loss of shape consistency, it still has the worst performance among the baselines. Besides, we can also see in Tab. II that OCFNet and ${}^{{\dagger}}$ still have the best performance. This experiment reveals the reason why Cam4DOcc suggests the inflated labels for GMO annotation in the occupancy forecasting task: Forecasting sophisticated future 3D structures of movable objects only using camera images is very difficult while forecasting inflated GMO potentially promotes more reliable and safer navigation in autonomous driving applications.
@@
${}^{{\dagger}}$ (ours) 29.84 17.72 23.78 25.53 17.81 21.67 26.53
[…]
Evaluation on forecasting fine-grained GMO, fine-grained GSO, and free space. In the fourth-level task, only OpenOccupancy-C and our OCFNet need to be retrained. As seen in Tab. IV, ${}^{{\dagger}}$ remains the best performance against all the other approaches on forecasting fine-grained objects of interest. Compared to the results in Tab. II, the GMO forecasting performance of OpenOccupancy-C and our OCFNet drops slightly due to additional artifacts introduced by the fine-grained GSO class.
@@ A Dataset Setup Details
We provide more details about our new dataset format for our Cam4DOcc benchmark by presenting statistics on the instance duration $[t_{in},t_{out}]$ after splitting the original nuScenes and Lyft-Level5 datasets to separate sequences mentioned in Sec. III-B. As shown in Fig. 6, most general movable objects (GMO) appear in at least two historical observations and all future observations ($[-2,4]$ and $[-1,4]$) in our benchmark. The long instance duration leads to an effective training strategy for the occupancy forecasting model. Besides, over 30% instances in the two datasets first appear in the current frame ($t=0$), which makes the model learn to forecast the object motion only according to their current location and surrounding conditions.
@@ C Study on Future Time Horizons
We further conduct a study on forecasting performance drops with different future time horizons. Since the occupancy grids of static objects do not change in the future time steps unless ground-truth annotations jitter, here we solely focus on the ability to forecast the future occupancy state of movable objects. In this experiment, we post the performance of OpenOccupancy-C, PowerBEV-3D, and our OCFNet for the first-level task and the second-level task since the baseline SPC fails to forecast the inflated GMO mentioned in Sec. V-B. As shown in Tab. VI, our ${}^{{\dagger}}$ remains the best performance for different time horizons in both tasks. In addition, all the baseline approaches show better performance on Lyft-Level5 than nuScenes as the time period for evaluating on Lyft-Level5 is relatively shorter. The closer the timestamp is to the current moment, the easier it is for all the baselines to forecast the occupancy status.
@@
where $TP_{t}$, $FP_{t}$, and $FN_{t}$ represent true positives, false positives, and false negatives at timestamp $t$. Note that in our work the predicted instance is regarded as one true positive once its IoU is greater than $0.2$ (adaptively chosen according to the level of IoU) and the corresponding instance ID is correctly tracked. The experimental results are shown in Tab. VII. Note that the instance prediction results of PowerBEV-3D are also from the duplication of forecasted 2D flow along the height dimension (same as its 3D extension of forecasted occupancy introduced in Sec. III-D). As can be seen, our proposed ${}^{{\dagger}}$ shows better 3D instance prediction ability than PowerBEV-3D on Lyft-Level5 while PowerBEV-3D outperforms our approach on nuScenes. In addition, ${}^{{\dagger}}$ improves the prediction of OCFNet by 30.2% and 13.7% on nuScenes and Lyft-Level5 respectively. The 2D-3D instance-based prediction baseline presents good instance prediction ability on nuScenes because 2D backward centripetal flow is easier to forecast than the 3D counterpart. On the contrary, our proposed method produces better forecasting results on Lyft-Level5, dominated by far better GMO occupancy forecasting quality of ${}^{{\dagger}}$ than that of PowerBEV-3D on this dataset. Therefore, in the 3D instance prediction task, we further propose a new baseline namely $*{}^{*}$, which combines the advantages of our original ${}^{{\dagger}}$ and PowerBEV-3D. The principle is that the 3D flow of the intersection GMO occupancy forecasted by the two methods follows PowerBEV-3D’s results, while the other GMO occupancy grids forecasted by ${}^{{\dagger}}$ have the motion flow generated by ${}^{{\dagger}}$ itself. Based on this setup, whether an occupancy grid is occupied totally depends on ${}^{{\dagger}}$, and its flow depends on the choice between ${}^{{\dagger}}$ and PowerBEV-3D. From Tab. VII, we can see that $*{}^{*}$ has the best 3D instance prediction performance, which enhances PowerBEV-3D by 6.7% on nuScenes and improves ${}^{{\dagger}}$ by 2.1% on Lyft-Level5.
@@ F Visualization of future GMO occupancy forecasted by OCFNet on Lyft-Level5
In this section, we present our proposed OCFNet forecasting inflated general movable objects of the Lyft-Level5 dataset. Fig. 10 and Fig. 11 show the results in small-scale and large-scale scenes respectively. The prediction results and ground truth from timestamps 1 to $N_{f}$ are assigned colors from dark to light. As to the small-scale scenes, the valid GMO over the future time horizon occupy relatively fewer volumes and both OCFNet and ${}^{{\dagger}}$ can capture their motion accurately. When it comes to the large-scale conditions, ${}^{{\dagger}}$ significantly outperforms OCFNet which only uses $\frac{1}{6}$ sequences for training. Therefore, when the driving scenario of the ego vehicle has few movable obstacles, such as in rural areas, OCFNet trained with limited data is enough to forecast the future occupancy of surrounding traffic participators. This can significantly improve the deployment efficiency of forecasting modules in autonomous driving systems by decreasing memory consumption and training period.

### v3

@@ III-A Task Definition
Given $N_{p}$ past and the current consecutive camera images $\mathcal{I}=\{I_{t}\}_{t=-N_{p}}^{0}$ as input, 4D occupancy forecasting aims to output the current occupancy $\mathbf{O}_{c}\in\mathbb{R}^{1\times H\times W\times L}$ and the future occupancy $\mathbf{O}_{f}\in\mathbb{R}^{N_{f}\times H\times W\times L}$ in a short time interval $N_{f}$, where $H$, $W$, $L$ represent the height, width, and length of the specific range defined in the present coordinate system ($t=0$). Each voxel of $\mathbf{O}_{f}$ has $N_{f}$ sequential states $\mathcal{S}=\{S_{t}\}_{t=1}^{N_{f}}$ to represent whether it is free or occupied in each future timestamp.
@@ III-B Dataset in New Format
Our Cam4DOcc benchmark introduces a new dataset format based on original nuScenes [29], nuScenes-Occupancy [13], and Lyft-Level5. As Fig. 2 illustrates, we first split the original nuScenes dataset into sequences with the time length of $N=N_{p}+N_{f}+1$. Then sequential semantic and instance annotations of movable objects are extracted for each sequence and collected into the GMO class, including bicycle, bus, car, construction, motorcycle, trailer, truck, and pedestrian. They are all transformed to the present coordinate system ($t=0$). After that, we voxelize the present 3D space and attach semantic/instance labels to the grids of movable objects using bounding boxes annotation. Notably, the invalid instance is discarded in this process once: (1) its visibility is under 40% over the 6 camera images if it is a newly appeared object in $N_{p}$ historical frames, (2) it first appears in $N_{f}$ incoming frames, or (3) it moves beyond the range ($H,W,L$) predefined at $t=0$. The visibility is quantified by the visible proportion of all pixels of the instance showing in camera images [29]. The sequential annotations are exploited to fill in missing intermediate instances based on constant velocity assumption [22, 44]. The same operations are also applied to the Lyft-Level5 dataset. The distribution of instance duration $t_{in},t_{out}]$ after the processing mentioned above is presented in supplementary Sec. A. Lastly, we generate 3D backward centripetal flow using the instance association in the annotations. Li et al. [22] introduced 2D backward centripetal flow to improve the efficiency of 2D instance prediction. Inspired by that, we calculate 3D backward centripetal flow pointing from the voxel at time $t$ to its corresponding 3D instance center at t1t-1. We exploit this 3D flow to improve the accuracy of camera-based 4D occupancy forecasting (see Sec. V-C).
@@ V-A Experimental Setups
Baseline setups. We choose the state-of-the-art camera-based approaches as the outset of each baseline proposed in Sec. III-D. For the static-world occupancy model, we use the camera baseline of OpenOccupancy [13] (OpenOccupancy-C) to estimate the occupancy state of the present frame, which is then regarded as the prediction of all the future time steps. For the voxelization of point cloud prediction, we use SurroundDepth [50] to estimate continuous surrounding depth maps, which are then downsampled to generate pseudo point clouds by ray casting. Based on sequential pseudo point clouds input, we then use PCPNet [37] to forecast incoming 3D point clouds, followed by Cylinder3D [46] to extract point-level GMO and GSO labels, and further voxelize the results into occupancy grids (SPC). For the 2D-3D instance-based prediction, we choose PowerBEV [22] to forecast occupancy semantics on BEV and then lift the 2D results to 3D space (PowerBEV-3D). As to our proposed OCFNet, we directly implement 4D occupancy forecasting end-to-end. Notably, PowerBEV is trained by the 2D ground-truth semantics and 2D flow projected to the BEV plane. Besides, only PowerBEV and OCFNet are trained with flow annotations from Cam4DOcc simultaneously since they both have the flow head. To show that our proposed OCFNet can generate good forecasted results even seeing limited training data, we report the performance of OCFNet only trained on $\frac{1}{6}$ training sequences as well as the performance of the one trained on all training sequences (OCFNet). OpenOccupancy-C, PowerBEV, and OCFNet are trained for 15 epochs using AdamW optimizer [51] with an initial learning rate 3e-4 and a weight decay of 0.01. SurroundDepth and Cylinder3D used in the point cloud prediction baseline are fine-tuned as their open sources suggest. PCPNet is firstly pretrained by range loss for 40 epochs using the same optimizer, but the initial learning rate is set to 1e-3. After that, it is further fine-tuned by Chamfer distance loss [52] for 10 epochs with a learning rate of 6e-4. All the networks mentioned above are trained with a batch size of 8 on 8 A100 GPUs. More details about the model parameters of our OCFNet are provided in supplementary Sec. B.
@@
Evaluation on forecasting inflated GMO. Results of the first task, forecasting inflated GMO on nuScenes and Lyft-Level5, are presented in Tab. I. Here, OpenOccupancy-C, PowerBEV, and OCFNet are trained only with inflated GMO labels, while PCPNet is trained by holistic point clouds. As shown, OCFNet and OCFNet outperform all other baselines, surpassing the BEV-based method by 12.4% and 13.3% in $\text{IoU}_{f}$ and $\tilde{\text{IoU}}_{f}$ on nuScenes. On Lyft-Level5, our OCFNet and OCFNet consistently outperforms PowerBEV-3D by 20.8% and 21.8% in $\text{IoU}_{f}$ and $\tilde{\text{IoU}}_{f}$. In addition, Fig. 5 shows the results of nuScenes GMO occupancy forecasted by our OCFNet and CFNet, which indicates that OCFNet trained only with limited data can still capture the motion of GMO occupancy grids reasonably. The visualization on Lyft-Level5 is shown in supplementary Sec. F. The baseline SPC cannot work well for the present frame and even tends to fail while forecasting future occupancy state. This is because movable objects are labeled as the inflated dense voxel grids in this task, while the voxelization of PCPNet outputs is from sparse point-level prediction. In addition, the shape of the predicted objects loses consistency significantly in future time steps. The performance of OpenOccupancy-C is much better than that of the point cloud prediction baseline but still has a weak ability to estimate present occupancy and forecast future occupancy compared to PowerEBV-3D and OCFNet.
@@
Evaluation on forecasting fine-grained GMO. We further report the occupancy estimation and forecasting performance on fine-grained general movable objects with nuScenes-Occupancy (the second-level task). In Tab. II, we exhibit how the IoU of the forecasted objects changes once the GMO annotations have fine-grained voxel format rather than the inflated one in the first-level task for training as well as evaluation. It can be seen that the IoU of GMO forecasted by all the methods except the point cloud prediction baseline decreases significantly because it is rather difficult to predict sophisticated moving 3D structures using past continuous camera images. In contrast, SPC presents slightly better performance compared to the results in Tab. I since the ground-truth labels are also fine-grained and sparser than the counterparts in the first-level task. However, due to the loss of shape consistency, it still has the worst performance among the baselines. Besides, we can also see in Tab. II that OCFNet and OCFNet still have the best performance. This experiment reveals the reason why Cam4DOcc suggests the inflated labels for GMO annotation in the occupancy forecasting task: Forecasting sophisticated future 3D structures of movable objects only using camera images is very difficult while forecasting inflated GMO potentially promotes more reliable and safer navigation in autonomous driving applications.
@@
OCFNet (ours) 29.84 17.72 23.78 25.53 17.81 21.67 26.53
Evaluation on forecasting inflated GMO, fine-grained GSO, and free space. Next, we compare the performance of different methods on forecasting inflated general movable objects, fine-grained general static objects, and free space (the third-level task). Here, we do not report the GSO results from the 2D-3D instance-based prediction since the fine-grained 3D structure of static foreground and background objects cannot be approximately estimated by lifting 2D voxel grids to 3D space. The experimental results are shown in Tab. III. SPC remains the worst in this experiment where the IoU of inflated GMO is consistent with the results of Tab. I. OCFNet and OCFNet outperform OpenOccupancy-C significantly in terms of estimating GMO occupancy in both present moment and future time steps. It also can be seen that by aggregating features of multiple past frames, OCFNet enhances the performance of GSO occupancy estimation of single-frame-based OpenOccupancy-C by 5.1% and 4.2% on IoUc and IoUf respectively. For OpenOccupancy-C and our OCFNet, the IoU values of future GSO slightly increase due to the jitter of ground truth annotations from nuScenes-Occupancy.
Evaluation on forecasting fine-grained GMO, fine-grained GSO, and free space. In the fourth-level task, only OpenOccupancy-C and our OCFNet need to be retrained. As seen in Tab. IV, OCFNet remains the best performance against all the other approaches on forecasting fine-grained objects of interest. Compared to the results in Tab. II, the GMO forecasting performance of OpenOccupancy-C and our OCFNet drops slightly due to additional artifacts introduced by the fine-grained GSO class.
@@ A Dataset Setup Details
We provide more details about our new dataset format for our Cam4DOcc benchmark by presenting statistics on the instance duration $t_{in},t_{out}]$ after splitting the original nuScenes and Lyft-Level5 datasets to separate sequences mentioned in Sec. III-B. As shown in Fig. 6, most general movable objects (GMO) appear in at least two historical observations and all future observations ([2,4][-2,4] and [1,4][-1,4]) in our benchmark. The long instance duration leads to an effective training strategy for the occupancy forecasting model. Besides, over 30% instances in the two datasets first appear in the current frame ($t=0$), which makes the model learn to forecast the object motion only according to their current location and surrounding conditions.
@@ C Study on Future Time Horizons
We further conduct a study on forecasting performance drops with different future time horizons. Since the occupancy grids of static objects do not change in the future time steps unless ground-truth annotations jitter, here we solely focus on the ability to forecast the future occupancy state of movable objects. In this experiment, we post the performance of OpenOccupancy-C, PowerBEV-3D, and our OCFNet for the first-level task and the second-level task since the baseline SPC fails to forecast the inflated GMO mentioned in Sec. V-B. As shown in Tab. VI, our OCFNet remains the best performance for different time horizons in both tasks. In addition, all the baseline approaches show better performance on Lyft-Level5 than nuScenes as the time period for evaluating on Lyft-Level5 is relatively shorter. The closer the timestamp is to the current moment, the easier it is for all the baselines to forecast the occupancy status.
@@
where $TP_{t}$, $FP_{t}$, and $FN_{t}$ represent true positives, false positives, and false negatives at timestamp $t$. Note that in our work the predicted instance is regarded as one true positive once its IoU is greater than $0.2$ (adaptively chosen according to the level of IoU) and the corresponding instance ID is correctly tracked. The experimental results are shown in Tab. VII. Note that the instance prediction results of PowerBEV-3D are also from the duplication of forecasted 2D flow along the height dimension (same as its 3D extension of forecasted occupancy introduced in Sec. III-D). As can be seen, our proposed OCFNet shows better 3D instance prediction ability than PowerBEV-3D on Lyft-Level5 while PowerBEV-3D outperforms our approach on nuScenes. In addition, OCFNet improves the prediction of OCFNet by 30.2% and 13.7% on nuScenes and Lyft-Level5 respectively. The 2D-3D instance-based prediction baseline presents good instance prediction ability on nuScenes because 2D backward centripetal flow is easier to forecast than the 3D counterpart. On the contrary, our proposed method produces better forecasting results on Lyft-Level5, dominated by far better GMO occupancy forecasting quality of OCFNet than that of PowerBEV-3D on this dataset. Therefore, in the 3D instance prediction task, we further propose a new baseline namely OCFNet, which combines the advantages of our original OCFNet and PowerBEV-3D. The principle is that the 3D flow of the intersection GMO occupancy forecasted by the two methods follows PowerBEV-3D’s results, while the other GMO occupancy grids forecasted by OCFNet have the motion flow generated by OCFNet itself. Based on this setup, whether an occupancy grid is occupied totally depends on OCFNet, and its flow depends on the choice between OCFNet and PowerBEV-3D. From Tab. VII, we can see that OCFNet has the best 3D instance prediction performance, which enhances PowerBEV-3D by 6.7% on nuScenes and improves OCFNet by 2.1% on Lyft-Level5.
@@ F Visualization of future GMO occupancy forecasted by OCFNet on Lyft-Level5
In this section, we present our proposed OCFNet forecasting inflated general movable objects of the Lyft-Level5 dataset. Fig. 10 and Fig. 11 show the results in small-scale and large-scale scenes respectively. The prediction results and ground truth from timestamps 1 to $N_{f}$ are assigned colors from dark to light. As to the small-scale scenes, the valid GMO over the future time horizon occupy relatively fewer volumes and both OCFNet and OCFNet can capture their motion accurately. When it comes to the large-scale conditions, OCFNet significantly outperforms OCFNet which only uses $\frac{1}{6}$ sequences for training. Therefore, when the driving scenario of the ego vehicle has few movable obstacles, such as in rural areas, OCFNet trained with limited data is enough to forecast the future occupancy of surrounding traffic participators. This can significantly improve the deployment efficiency of forecasting modules in autonomous driving systems by decreasing memory consumption and training period.